# CAD-VAE: Leveraging Correlation-Aware Latents for Comprehensive Fair Disentanglement

### v1

@@ 3.3 Direct Minimization of Conditional Mutual Information
where $p_{\theta}\bigl{(}\hat{y}\mid z_{S}^{(i)}\bigr{)}\!=\!f_{y\_op}\bigl{(}z_{S}^{(i)}\bigr{)}.$ Here, $z_{S}^{(i)}$ denotes the $z_{S}$ sample from the $i$-th element in a mini-batch of size $|B|$; the distribution $q_{\phi}(z_{S}\mid x)$ is given by the encoder. Similar calculation to $H_{\phi}(\hat{S}\mid z_{Y})$, see Appendix C for completed calculation formula. During this optimization process, the opponent classifier parameters $\omega_{y\_op}$ and $\omega_{s\_op}$ are frozen.
3.4 Learning Relevance Between Target And Sensitive Information
@@
$H_{\phi}(*\mid*)$ stands for the conditional entropy. Maximizing $H_{\phi}(\hat{Y}\mid z_{R})$, $H_{\phi}(\hat{S}\mid z_{R})$ avoid $z_{R}$ capture all information of $Y$ and $S$: $Y\cap S$, which will lead to the Information Bottleneck phenomenon[15, 6, 19] i.e $z_{R}$ capture all the information about target attribute $Y$ and sensitive attribute $S$, degenerating disentanglement performance. On the other hand, minimizing $H_{\phi}(\hat{Y}\mid Y,z_{R})$ and $H_{\phi}(\hat{S}\mid S,z_{R})$ encourage $z_{R}$ capture relevant information between $Y$ and $S$, which is achieved by marginalizing latent codes in different attribute groups as shown in next.
@@
The calculation of $H_{\phi}(\hat{S}\mid z_{R})$ and $H_{\phi}(\hat{S}\mid S,z_{R})$ are similar to $H_{\phi}(\hat{Y}\mid z_{R})$, $H_{\phi}(\hat{Y}\mid Y,z_{R})$ respectively, see Appendix C for completed calculation formula. Plugging these estimates(21)(LABEL:cal_H_Y_Y_zr) back into (17), shared feature between $Y$ and $S$ will be learned in $z_{R}$ while getting rid of Information Bottleneck phenomenon. Note that the classifier parameters $\omega_{y}$ and $\omega_{s}$ are remain freeze when (17) optimizing.
3.5 Final Objective Function
@@
Specifically, in (24), we jointly update $(\theta,\phi,\omega_{y},\omega_{s})$ by minimizing the VAE loss (2) alongside the main classification losses (3) and (4), which together reformulate the ELBO. We further include the CMI loss (15) to reduce unwanted information leakage, the LRI loss (17) to capture shared patterns in $z_{R}$, and the TC penalty (1) to promote factorization among the latent codes $(z_{Y},z_{R},z_{S})$.
@@
The hyperparameters $\lambda_{CMI},\lambda_{LRI}>0$ control the relative importance of these terms, ensuring each network component learns its designated function while enforcing minimal information leakage, preserving shared information in $z_{R}$ and maintaining the salient factors for $Y$ and $S$ in $z_{Y}$ and $z_{S}$ respectively. See hypermeter analysis in Appendix D.
4 Experiment
Figure 3: Examples of Fair Counterfactual Generation. The first row shows the source and reference images. Rows 2–5 display counterfactuals obtained by replacing latent subspaces $z_{X}$, $z_{Y}$, $z_{S}$, and $[z_{S},z_{R}]$, respectively. Notably, the replacement with $[z_{S},z_{R}]$ (row 5) naturally adapts sensitive features for different sensitive attributes without domain knowledge. (mustache for men and makeup for women).
@@ 4 Experiment
To ensure a rigorous and comprehensive evaluation, we conduct experiments comparing our proposed method with a diverse set of state-of-the-art approaches across multiple categories of learning paradigms in various tasks. Specifically, we include FairFactorVAE [28], FairDisCo [27], FFVAE [6], GVAE [7], ODVAE [42] and FADES [15], shown in Section 2.1. As for traditional correlation-aware learning that is discussed in Section 2.2, since it require additional annotated data to build causal graph, we except them in our experiment. Further details on the experimental setup and additional discussions are provided in Appendix D.
4.1 Fair Classification
                 Downstream Classification Performance Methods CelebA [29] UTKFace [49] Dogs and Cats [37] Color bias MNIST [17] Acc $\uparrow$ EOD $\downarrow$ DP $\downarrow$ Acc $\uparrow$ EOD $\downarrow$ DP $\downarrow$ Acc $\uparrow$ EOD $\downarrow$ DP $\downarrow$ Acc $\uparrow$ EOD $\downarrow$ DP $\downarrow$   FADES [15] [CVPR’24] 0.918 0.034 0.135 0.812 0.059 0.139 0.769 0.058 0.086 0.973 0.094 0.160 GVAE [7] [CVPR’20] 0.919 0.047 0.131 0.819 0.204 0.197 0.748 0.064 0.131 0.961 0.109 0.176 FFVAE [6] [PMLR’19] 0.892 0.076 0.072 0.766 0.269 0.201 0.729 0.059 0.110 0.952 0.081 0.092 ODVAE [42] [ECCV’20] 0.886 0.039 0.103 0.736 0.165 0.210 0.689 0.051 0.038 0.957 0.247 0.162 FairDisCo [27] [KDD’22] 0.839 0.074 0.051 0.766 0.266 0.200 0.680 0.115 0.111 0.949 0.129 0.136 FairFactorVAE [28] 0.914 0.055 0.136 0.720 0.096 0.134 0.707 0.055 0.110 0.957 0.096 0.128 CAD-VAE (Ours) 0.923 0.021 0.112 0.817 0.045 0.137 0.773 0.048 0.090 0.977 0.076 0.118
@@ 4.1 Fair Classification
The objective of fair classification is to achieve a balance between minimizing fairness violations and maintaining high predictive performance. To evaluate the effectiveness of our proposed method, we conduct experiments on a diverse set of benchmark fairness datasets. For facial attribute classification tasks, we utilize the CelebA [29] and UTKFace [49] datasets. Following prior works [44, 47, 48, 15], we set the CelebA classification task to predict the “Smiling” attribute, while for UTKFace, the objective is to classify whether a person depicted in the image is over 35 years old, with gender serving as the sensitive attribute. Additionally, the Dogs and Cats dataset [37] is used to distinguish between dogs and cats, with fur color as the sensitive attribute. Furthermore, we assess fair classification performance using the Colored MNIST dataset [18, 17, 35], which incorporates a controlled color bias in the standard MNIST dataset to simulate spurious correlations. To assess fairness violations, we use standard metrics including Demographic Parity (DP) [2] and Equalized Odds (EOD) [10]. See detailed experimental setup in Appendix D.
The result of fair classification can be seen in Table 1. Across all evaluated datasets, our method consistently achieves state-of-the-art classification accuracy and fairness, validating its effectiveness in robust disentanglement by preserving high-quality target-related information while minimizing sensitive attribute leakage. Among the baselines, FADES achieves competitive accuracy and strong fairness. However, its inability to strictly minimize conditional mutual information results in sensitive information leakage and an excessive capture of target-related data in the relevance latent code $z_{R}$, which limits classification performance. In contrast, our approach directly minimizes the conditional mutual information between $z_{Y}$ and $z_{S}$ regarding their respective opposite attributes $S$ and $Y$, and incorporates an explicit relevance learning strategy, as detailed in Section 3.3 and Section 3.4. This design ensures that $z_{R}$ captures sufficient, but not excessive, information about $Y$ and $S$, thereby enhancing fairness while preserving essential target information. More fair classification experiment results can be seen in Appendix E.
4.2 Fair Counterfactual Generation
We evaluate our approach on the CelebA dataset [29], a widely-used benchmark for facial attribute manipulation. We select Smiling as the target label $Y$ and Gender as the sensitive attribute $S$. In our experiments, we substitute specific latent code of source images with reference images, including $z_{X}$, $z_{Y}$, $z_{S}$, and $[z_{S},z_{R}]$. Figure 3 illustrates the generated counterfactuals, with the first row showing source and reference images and subsequent rows demonstrating the effects of substituting each latent code.
@@ 4.2 Fair Counterfactual Generation
The experimental results demonstrate the effectiveness of our method in generating fair counterfactuals. As shown in Figure 3, substituting $[z_{S},z_{R}]$ (Row 5) leads to a natural adaptation of sensitive-relevant features without domain-specific knowledge. For instance, the model automatically adds makeup to female images and a mustache to male images, highlighting the semantic alignment of $z_{R}$ with both the target and sensitive attributes. Compared to substituting only $z_{S}$, our approach achieves more interpretable translation, ensuring that fairness is maintained throughout the counterfactual generation. More fair counterfactual generation experiment results can be seen in Appendix E.
@@
To quantitatively assess the quality of the generated counterfactuals, we compare evaluation metrics between the direct reconstruction of the input image and the reconstructions obtained by randomly permuting $z_{Y}$ and $z_{S}$ within the evaluation set. Specifically, we use the FID [12, 15] to assess reconstruction fidelity and the Inception Score (IS) [5] to evaluate semantic and perceptual quality. Lower $\Delta\text{FID}$ values indicate minimal distortion and higher translation quality, while lower $\Delta\text{IS}$ values suggest that semantic and perceptual attributes are well preserved. Detailed experimental settings are provided in Appendix D.
@@
4.3 Fair Fine-Grained Image Editing
@@
In our experiments, following the setup in Section 4.2 where Smiling is the target label $Y$ and Gender is the sensitive attribute $S$, we generate interpolated latent codes between source and reference images. In the blue-framed subfigure of Figure 1, images are generated by interpolating $z_{Y}$ and $z_{S}$: the horizontal axis shows the transition of $z_{S}$ from the source to the reference image, while the vertical axis shows the corresponding change in $z_{Y}$. During this interpolation, $z_{R}$ and $z_{X}$ remain unchanged, with the interpolation parameters for both $z_{Y}$ and $z_{S}$ set to $\lambda\in\{0,0.33,0.66,1\}$.
@@
Figure 1 demonstrates a smooth transformation of each attribute, with modifications in one latent code minimally affecting the others, a key characteristic of effective disentanglement. Specifically, as the correlated latent code $z_{R}$ captures sensitive relevant information, we can explicitly control these properties: in the left subfigure, we gradually introduce makeup (such as enhanced lipstick and eyeshadow), while in the right subfigure, we progressively add a mustache. More fair fine-grained image editing experiment results can be seen in Appendix E.
Similarly, we measure $\Delta\text{FID}$ and $\Delta\text{IS}$ to quantitatively assess the quality of fine-grained image editing. Unlike the evaluation setup in Section 4.2, we compute the differences in evaluation metrics between the direct reconstruction of the input image and the reconstructions obtained through latent code traversals for each $\lambda$ combination. Detailed experimental settings are provided in Appendix D. Table 3 summarizes the comparison results, showing that our method exhibits both lower $\Delta\text{FID}$ and $\Delta\text{IS}$ values compared to other fair generation methods. These results indicate that our fine-grained image editing approach not only ensures smoother attribute transformations and superior image fidelity, but also allows for more precise control of task-relevant features.
4.4 Fair Text-to-Image Editing
@@
To further validate the capability and explore the applicability of our method, we integrated it as an adaptor on top of a pre-trained, frozen CLIP image encoder [39] and trained it on Facet dataset [9] to enhance fairness in vision-language tasks. Table 4 presents the experimental results. These results demonstrate that our approach significantly improves fairness without compromising performance compared to the linear probing baseline (ERM), underscoring its potential for a range of vision-language tasks with fairness considerations.
Furthermore, we applied our method in StyleCLIP [38] as a fair discriminator to address inherent fairness issues, such as career-gender biases, which persist even when an identity preservation loss is employed. As illustrated in Figure 8, StyleCLIP [38] exhibits a bias by correlating the role of “dancer” with a specific gender. In contrast, our method effectively mitigates this bias while maintaining the efficacy of attribute modification. See Appendix E for details.

### v2

@@ Introduction
where $p_{\theta}\bigl(\hat{y}\mid z_{S}^{(i)}\bigr)\!=\!f_{y\_op}\bigl(z_{S}^{(i)}\bigr).$ Here, $z_{S}^{(i)}$ denotes the $z_{S}$ sample from the $i$-th element in a mini-batch of size $|B|$; the distribution $q_{\phi}(z_{S}\mid x)$ is given by the encoder. Similar calculation to $H_{\phi}(\hat{S}\mid z_{Y})$, see Appendix 3 for completed calculation formula. During this optimization process, the opponent classifier parameters $\omega_{y\_op}$ and $\omega_{s\_op}$ are frozen.
                 Downstream Classification Performance Methods CelebA (Liu et al. 2015) UTKFace (Zhang, Song, and Qi 2017) Dogs and Cats (Parkhi et al. 2012) Color bias MNIST (Kim et al. 2019) Acc $\uparrow$ EOD $\downarrow$ DP $\downarrow$ Acc $\uparrow$ EOD $\downarrow$ DP $\downarrow$ Acc $\uparrow$ EOD $\downarrow$ DP $\downarrow$ Acc $\uparrow$ EOD $\downarrow$ DP $\downarrow$   FADES (Jang and Wang 2024) [CVPR’24] 0.918 0.034 0.135 0.812 0.059 0.139 0.769 0.058 0.086 0.973 0.094 0.160 GVAE (Ding et al. 2020) [CVPR’20] 0.919 0.047 0.131 0.819 0.204 0.197 0.748 0.064 0.131 0.961 0.109 0.176 FFVAE (Creager et al. 2019) [PMLR’19] 0.892 0.076 0.072 0.766 0.269 0.201 0.729 0.059 0.110 0.952 0.081 0.092 ODVAE (Sarhan et al. 2020) [ECCV’20] 0.886 0.039 0.103 0.736 0.165 0.210 0.689 0.051 0.038 0.957 0.247 0.162 FairDisCo (Liu et al. 2022) [KDD’22] 0.839 0.074 0.051 0.766 0.266 0.200 0.680 0.115 0.111 0.949 0.129 0.136 FairFactorVAE (Liu, Sun, and Zhao 2023) 0.914 0.055 0.136 0.720 0.096 0.134 0.707 0.055 0.110 0.957 0.096 0.128 CAD-VAE (Ours) 0.939 0.021 0.065 0.828 0.045 0.137 0.781 0.048 0.069 0.984 0.076 0.108
Learning Relevance Between Target And Sensitive Information
@@
$H_{\phi}(*\mid*)$ stands for the conditional entropy. Maximizing $H_{\phi}(\hat{Y}\mid z_{R})$, $H_{\phi}(\hat{S}\mid z_{R})$ avoid $z_{R}$ capture all information of $Y$ or $S$ solely, which will lead to the Information Bottleneck phenomenon(Jang and Wang 2024; Creager et al. 2019; Kim and Mnih 2018) i.e $z_{R}$ capture all the information about target attribute $Y$ and sensitive attribute $S$: $Y\cup S$, degenerating disentanglement performance. On the other hand, minimizing $H_{\phi}(\hat{Y}\mid Y,z_{R})$ and $H_{\phi}(\hat{S}\mid S,z_{R})$ enforce $z_{R}$ determine $\hat{Y}$ or $\hat{S}$ only within each $Y$ or $S$ subgroup, so that encourage $z_{R}$ capture information both relevant to $Y$ and $S$: $Y\cap S$.
In contrast to FADES(Jang and Wang 2024), which exhibits a conflict between the disentanglement term and the regularization term, our method achieves orthogonality between CMI (20) and LRI (22), ensuring fair disentanglement while preserving robust representations. See Appendix 5 for details.
@@
The calculation of $H_{\phi}(\hat{S}\mid z_{R})$ and $H_{\phi}(\hat{S}\mid S,z_{R})$ are similar to $H_{\phi}(\hat{Y}\mid z_{R})$, $H_{\phi}(\hat{Y}\mid Y,z_{R})$ respectively, see Appendix 3 for completed calculation formula. Plugging these estimates(26)(LABEL:cal_H_Y_Y_zr) back into (22), shared feature between $Y$ and $S$ will be learned in $z_{R}$ while getting rid of Information Bottleneck phenomenon. Note that the classifier parameters $\omega_{y}$ and $\omega_{s}$ remain frozen when optimizing (22).
Final Objective Function
@@
Specifically, in (29), we jointly update $(\theta,\phi,\omega_{y},\omega_{s})$ by minimizing the VAE loss (LABEL:vae_loss) alongside the main classification losses (8) and (9), which together reformulate the ELBO. We further include the CMI loss (20) to reduce unwanted information leakage, the LRI loss (22) to capture shared patterns in $z_{R}$, and the TC penalty (6) to promote factorization among the latent codes $(z_{Y},z_{R},z_{S})$.
@@
The hyperparameters $\lambda_{CMI},\lambda_{LRI}>0$ control the relative importance of these terms, ensuring each network component learns its designated function while enforcing minimal information leakage, preserving shared information in $z_{R}$ and maintaining the salient factors for $Y$ and $S$ in $z_{Y}$ and $z_{S}$ respectively. See hypermeter analysis in Appendix 6.
Experiment
@@
To ensure a rigorous and comprehensive evaluation, we conduct experiments comparing our proposed method with a diverse set of state-of-the-art approaches across multiple categories of learning paradigms in various tasks. Specifically, we include FairFactorVAE (Liu, Sun, and Zhao 2023), FairDisCo (Liu et al. 2022), FFVAE (Creager et al. 2019), GVAE (Ding et al. 2020), ODVAE (Sarhan et al. 2020) and FADES (Jang and Wang 2024), shown in Section 2.1. As for traditional correlation-aware learning that is discussed in Section 2.2, since it require additional annotated data to build causal graph, we except them in our experiment.
Fair Classification
@@
The objective of fair classification is to achieve a balance between minimizing fairness violations and maintaining high predictive performance. To evaluate the effectiveness of our proposed method, we conduct experiments on a diverse set of benchmark fairness datasets. For facial attribute classification tasks, we utilize the CelebA (Liu et al. 2015) and UTKFace (Zhang, Song, and Qi 2017) datasets. Following prior works (Wang et al. 2022; Xu et al. 2020; Zeng et al. 2022; Jang and Wang 2024), we set the CelebA classification task to predict the “Smiling” attribute, while for UTKFace, the objective is to classify whether a person depicted in the image is over 35 years old, with gender serving as the sensitive attribute. Additionally, the Dogs and Cats dataset (Parkhi et al. 2012) is used to distinguish between dogs and cats, with fur color as the sensitive attribute. Furthermore, we assess fair classification performance using the Colored MNIST dataset (Kim, Lee, and Choo 2021; Kim et al. 2019; Nam et al. 2020), which incorporates a controlled color bias in the standard MNIST dataset to simulate spurious correlations. To assess fairness violations, we use standard metrics including Demographic Parity (DP) (Barocas and Selbst 2016) and Equalized Odds (EOD) (Hardt, Price, and Srebro 2016). See detailed experimental setup in Appendix 6. The result of fair classification can be seen in Table 1. Across all evaluated datasets, our method consistently achieves state-of-the-art classification accuracy and fairness, validating its effectiveness in robust disentanglement by preserving high-quality target-related information while minimizing sensitive attribute leakage.
Table 2: Fair Classification on 95% Color Bias MNIST
  Metric Ours FADES GVAE FFVAE ODVAE FairFactorVAE Accuracy $\uparrow$ 0.867 0.782 0.771 0.744 0.721 0.807 EOD $\downarrow$ 0.141 0.174 0.244 0.190 0.210 0.195 DP $\downarrow$ 0.167 0.201 0.265 0.213 0.262 0.221
To assess the robustness and effectiveness of these methods, we further conduct classification experiments under an extreme imbalance bias setting, which is commonly encountered in practical applications. In the MNIST experiment, we set the color bias rate to 95% to simulate a strong correlation between the target attribute and the sensitive attribute. The results, shown in Table 2, demonstrate that our method outperforms existing approaches in both disentanglement ability and robust representation preservation.
Figure 2: Examples of Fair Counterfactual Generation. Zoom in to check. The first row shows the source and reference images. Rows 2–5 display counterfactuals obtained by replacing latent subspaces $z_{X}$, $z_{Y}$, $z_{S}$, and $z_{S},z_{R}]$, respectively. Notably, the replacement with $z_{S},z_{R}]$ (row 5) naturally adapts sensitive features for different sensitive attributes without domain knowledge. (mustache for men and makeup for women).
Fair Counterfactual Generation
@@
We evaluate our approach on the CelebA dataset (Liu et al. 2015), a widely-used benchmark for facial attribute manipulation. We select Smiling as the target label $Y$ and Gender as the sensitive attribute $S$. In our experiments, we substitute specific latent code of source images with reference images, including $z_{X}$, $z_{Y}$, $z_{S}$, and $z_{S},z_{R}]$. Figure 2 illustrates the generated counterfactuals, with the first row showing source and reference images and subsequent rows demonstrating the effects of substituting each latent code. The experimental results demonstrate the effectiveness of our method in generating fair counterfactuals. As shown in Figure 2, substituting $z_{S},z_{R}]$ (Row 5) leads to a natural adaptation of sensitive-relevant features without domain-specific knowledge. For instance, the model automatically adds makeup to female images and a mustache to male images, highlighting the semantic alignment of $z_{R}$ with both the target and sensitive attributes. Compared to substituting only $z_{S}$, our approach achieves more interpretable translation, ensuring that fairness is maintained throughout the counterfactual generation. More fair counterfactual generation experiment results can be seen in Appendix 7.
@@
To quantitatively assess the quality of the generated counterfactuals, we compare evaluation metrics between the direct reconstruction of the input image and the reconstructions obtained by randomly permuting $z_{Y}$ and $z_{S}$ within the evaluation set. Specifically, we use the FID (Heusel et al. 2017; Jang and Wang 2024) to assess reconstruction fidelity and the Inception Score (IS) (Chong and Forsyth 2020) to evaluate semantic and perceptual quality. Lower $\Delta\text{FID}$ values indicate minimal distortion and higher translation quality, while lower $\Delta\text{IS}$ values suggest that semantic and perceptual attributes are well preserved. Detailed experimental settings are provided in Appendix 6.
Figure 3: Examples of Fair Fine-Grained Image Editing. Zoom in to check. The leftmost column shows the source and reference images. The blue-framed section displays images generated by interpolating $z_{Y}$ and $z_{S}$ (with $z_{R}$ and $z_{X}$ fixed), where the horizontal axis varies $z_{S}$ and the vertical axis varies $z_{Y}$. The red-framed section illustrates images produced by interpolating $z_{Y}$ and $z_{R}$ (with $z_{S}$ fully replaced by the reference and $z_{X}$ constant). Modification in one latent code minimally affecting others, harness $z_{R}$ to edit sensitive relevant feature(makeup or mustache).
@@
Fair Fine-Grained Image Editing
@@
In our experiments, following the setup in the previous section where Smiling is the target label $Y$ and Gender is the sensitive attribute $S$, we generate interpolated latent codes between source and reference images. In the blue-framed subfigure of Figure 6, images are generated by interpolating $z_{Y}$ and $z_{S}$: the horizontal axis shows the transition of $z_{S}$ from the source to the reference image, while the vertical axis shows the corresponding change in $z_{Y}$. During this interpolation, $z_{R}$ and $z_{X}$ remain unchanged, with the interpolation parameters for both $z_{Y}$ and $z_{S}$ set to $\lambda\in\{0,0.33,0.66,1\}$.
@@
Figure 6 demonstrates a smooth transformation of each attribute, with modifications in one latent code minimally affecting the others, a key characteristic of effective disentanglement. Specifically, as the correlated latent code $z_{R}$ captures sensitive relevant information, we can explicitly control these properties: in the left subfigure, we gradually introduce makeup (such as enhanced lipstick and eyeshadow), while in the right subfigure, we progressively add a mustache. More fair fine-grained image editing experiment results can be seen in Appendix 7.
Similarly, we measure $\Delta\text{FID}$ and $\Delta\text{IS}$ to quantitatively assess the quality of fine-grained image editing. Unlike the evaluation setup in the previous section, we compute the differences in evaluation metrics between the direct reconstruction of the input image and the reconstructions obtained through latent code traversals for each $\lambda$ combination. Detailed experimental settings are provided in Appendix 6. Table 4 summarizes the comparison results, showing that our method exhibits both lower $\Delta\text{FID}$ and $\Delta\text{IS}$ values compared to other fair generation methods. These results indicate that our fine-grained image editing approach not only ensures smoother attribute transformations and superior image fidelity, but also allows for more precise control of task-relevant features.
Fair Text-to-Image Editing
To further validate the capability and explore the applicability of our method, we integrated it as an adaptor on top of a pre-trained, frozen CLIP image encoder (Radford et al. 2021; Xiao et al. 2025c, a, b) and trained it on Facet dataset (Gustafson et al. 2023) to enhance fairness in vision-language tasks. Table 5 presents the experimental results. These results demonstrate that our approach significantly improves fairness without compromising performance compared to the linear probing baseline (ERM), underscoring its potential for a range of vision-language tasks with fairness considerations.
Furthermore, we applied our method in StyleCLIP (Patashnik et al. 2021) as a fair discriminator to address inherent fairness issues, such as career-gender biases, which persist even when an identity preservation loss is employed. As illustrated in Figure 8, StyleCLIP (Patashnik et al. 2021) exhibits a bias by correlating the role of “dancer” with a specific gender. In contrast, our method effectively mitigates this bias while maintaining the efficacy of attribute modification. See Appendix 7 for details.